INFRARED PHOTODISSOCIATION SPECTROSCOPY OF MASS-SELECTED METAL ION-BENZENE COMPLEXES

\underline{M. A. DUNCAN}, DEPARTMENT OF CHEMISTRY, UNIVERSITY OF GEORGIA, ATHENS, GA. 30602; D. van HEIJNSBERGEN, G. von HELDEN, and G. MEIJER, F. O. M. LABORA-TORY FOR PLASMAPHYSICS RIJNHUISEN, EDINSONBAAN 14, NL-3439 MN NIEUWEGEIN, THE NETHERLANDS.

Metal ion-benzene complexes of the form M^{+}(benzene)_{1,2} (M=V, Ti, Al) are produced by laser vaporization in a pulsed nozzle source, mass-selected in an ion-trap mass spectrometer and excited in the infrared with a free electron laser. Multiphoton dissociation occurs by the elimination of benzene molecules. The photofragment yield versus wavelength produces IR resonance-enhanced multiphoton photodissociation (IR-REMPD) spectroscopy for these complexes. Vibrational bands in the 600-1800cm^{−1}region are characteristic of the benzene molecular moiety perturbed by the metal bonding. Experimental data are compared to the IR spectra derived from density functional calculations. Vibrational patterns in V^{+}(C_{6}H_{6}) indicate that the metal is bound in an η^{6} π-bonding configuration, while V^{+}(C_{6}H_{6})_{2} is a sandwich. Trapped-ion IR-REMPD is a general method to access the vibrational spectra of novel organometallic complexes.